# Equivariant Hilbert and Ehrhart series under translative group actions

### v2

@@ 4.1. The Cohen–Macaulay approach
Let $\mathcal{O}\in\Sigma/G$. Each $\rho(g)$ permutes the faces of $\Sigma$ belonging to the orbit $\mathcal{O}$, and one can thus define the character $\mathfrak{f}_{\mathcal{O}}^{\Sigma}$ of the associated complex permutation representation. (When $\mathcal{O}$ is the orbit of the empty face of $\Sigma$, we set $\mathfrak{f}_{\mathcal{O}}^{\Sigma}=1$.) Defining $\mathfrak{f}_{i-1}^{\Sigma}$ to be the character of the complex permutation representation on the set of all $(i-1)$-dimensional faces of $\Sigma$, it follows that $\mathfrak{f}_{i-1}^{\Sigma}=\sum_{\mathrm{rk}(\mathcal{O})=i}\mathfrak{f}^{\Sigma}_{\mathcal{O}}$.
@@
where we are setting $10^{0}=$ 1 by convention. Since the coloring $\gamma$ is preserved by the group action and the action permutes the vertices of $\Sigma$, we have that
@@
for some $a^{(i)}_{0},\ldots a^{(i)}_{c}\in\mathbb{K}$. Now define a coloring of $\Sigma$ by assigning to every vertex $v$ the unique color $\gamma(v)\in\{0,\ldots,c\}$ such that $v\in Gw_{\gamma(v)}$. This coloring is obviously preserved by $G$. In order to prove that the action of $G$ is translative, it is enough to prove that the coloring $\gamma$ is proper. If $\Sigma$ is 00-dimensional, this is trivially true. Assume thus that $\Sigma$ has dimension at least $1$ and consider any nonempty edge $\sigma=\{v,w\}\in\Sigma$. By the Kind–Kleinschmidt criterion (Lemma 4.4), the vector subspace $\langle\theta_{0}|_{\sigma},\theta_{1}|_{\sigma},\ldots\theta_{d}|_{\sigma}\rangle$ has dimension $2$. By construction, the coefficient of $x_{v}$ (respectively, $x_{w}$) in $\theta_{i}$ is $a^{(i)}_{\gamma(v)}$ (respectively, $a^{(i)}_{\gamma(w)}$). As a consequence, the $2$-dimensional subspace $\langle\theta_{0}|_{\sigma},\theta_{1}|_{\sigma},\ldots\theta_{d}|_{\sigma}\rangle$ is the image of the $2\times(d+1)$ matrix
@@
The proof of Proposition 4.6 shows that the statement holds also for those finite fields $\mathbb{K}$ containing at least as many distinct elements as the number of vertices of $\Sigma$.
@@
For a simple example where Theorem 4.8 can be used, we refer the reader to Example 3.4. Indeed, the ring $R$ in Example 3.4 is the Stanley–Reisner ring over $\mathbb{C}$ of the 00-dimensional (and hence Cohen–Macaulay) simplicial complex $\Sigma$ having two points as its facets.
@@ 4.2. The colorful approach
Let us write $\mathfrak{f}^{\Sigma}_{S}$ for the character of the associated complex permutation representation. Note that $\mathfrak{f}^{\Sigma}_{S}$ is also the character of the permutation representation of $G$ on the standard basis of the vector space $C_{|S|-1}(\Sigma|_{S},\mathbb{C})$ of $(|S|-1)$-chains of $\Sigma|_{S}$ constructed with the orientation chosen as in Remark 4.10.
@@
Let $S\subseteq\Gamma$. By [Sta79, Theorem 4.3], the $(|S|-1)$-dimensional color-selected subcomplex $\Sigma|_{S}$ is Cohen–Macaulay, and hence $\widetilde{H}_{i}(\Sigma|_{S},\mathbb{C})$ vanishes for every $i<|S|-1$. Then $\mathfrak{h}_{S}^{\Sigma}$ is the character of the $G$-representation on $\widetilde{H}_{|S|-1}(\Sigma|_{S},\mathbb{C})$ from Lemma 4.12. ∎
@@ 5.1. The equivariant Ehrhart series of a lattice polytope
Now let $\widetilde{M}:=M\oplus\mathbb{Z}\simeq\mathbb{Z}^{d+1}$ and, for every $\phi\in\operatorname{Aff}(M)$ and every $x\oplus u\in\widetilde{M}$, let $\widetilde{\phi}(x\oplus u):=\phi(x)\oplus u$. Then $\widetilde{\phi}\in\mathrm{GL}(\widetilde{M})$, and so every affine representation $\rho$ defines a linear representation $\widetilde{\rho}$. Tensoring with $\mathbb{C}$, any affine representation $\rho:G\to\operatorname{Aff}(\mathbb{C}^{d})$ determines a (linear) representation $\widetilde{\rho}:G\to\operatorname{GL}_{d+1}(\mathbb{C})$ of $G$ on $\widetilde{M}\otimes\mathbb{C}\simeq\mathbb{C}^{d+1}$ whose image is in the subgroup of transformations preserving the affine subspace $\mathbb{C}^{d}\times\{1\}\subseteq\widetilde{M}\otimes\mathbb{C}$, see [Sta23, Section 2.3].
@@
where $h^{*}(P,\rho;t)$ is a power series in $R_{G}[[t]]$ and not necessarily a polynomial! Note that, when all coefficients of $h^{*}(P,\rho;t)$ are effective, the power series $h^{*}(P,\rho;t)$ is actually a polynomial, as follows from considering its evaluation at $\mathrm{id}_{G}$: see e.g. [EKS23, Lemma 2.8]. It is currently open whether polynomiality and effectiveness of $h^{*}(P,\rho;t)$ are equivalent conditions; this is conjectured to be true in [Sta23, Conjecture 1.2].
@@ 6. Order polytopes
In the setting of Remark 6.5, note that if $S\subseteq\{0,1,\ldots,r\}$ contains at least one of 00 and $r$, then $\Delta({Q})|_{S}$ contains at least one of the cone points $\hat{0}$ and $\hat{1}$ and is hence contractible. This implies that its reduced homology vanishes and, thus, $\mathfrak{h}_{S}^{\Delta(Q)}=\mathfrak{b}_{S}^{\Delta(Q)}=0$.
@@
By Lemma 6.4, the action $\rho^{*}$ on $\mathscr{T}_{X}$ is translative and preserves a balanced coloring, while $\rho$ describes the action on $\Delta(J(X))$ induced by the equivariant isomorphism of abstract simplicial complexes $\phi\colon\underline{\mathscr{T}_{X}}\simeq\Delta(J(X))$. To compute the equivariant Ehrhart series of $\mathscr{O}(X)$ with respect to $\rho^{*}$, by Theorem 5.2 it is enough to compute the equivariant Hilbert series of the abstract simplicial complex $\underline{\mathscr{T}_{X}}$ with respect to $\rho^{*}$. By the equivariant isomorphism of abstract simplicial complexes $\phi\colon\underline{\mathscr{T}_{X}}\simeq\Delta(J(X))$ from Lemma 6.4, this is in turn equal to the equivariant Hilbert series of the abstract simplicial complex $\Delta(J(X))$ with respect to $\rho$. The poset $J(X)$ is bounded with $\hat{0}=\emptyset$ and $\hat{1}=X$, and it is ranked by cardinality of its elements. By Remark 6.6, we can restrict to color sets $S\subseteq[d-1]$. The claim about the equivariant flag $h$-vector refining the equivariant $h$-vector follows from Proposition 4.14.(ii). ∎
@@
Let $H_{d}$ denote the poset consisting of a single antichain with $d$ elements. Then $\mathscr{O}(H_{d})$ is the unit cube $[0,1]^{d}$. Consider the group $G=S_{d}$, the symmetric group on $d$ elements, acting on $H_{d}$. The poset $J(H_{d})$ is the boolean poset $B_{d}$ of all subsets of $[d]$, and the induced action of $S_{d}$ is the natural action permuting elements of $[d]$. As in Theorem 6.7, let $\rho$ denote the corresponding action on the Stanley-Reisner ring of $\Delta(J(H_{d}))=\Delta(B_{d})$ and let $\rho^{*}$ denote the representation of $S_{d}$ permuting coordinates in $\mathbb{R}^{H_{d}}$. Then $\rho^{*}$ preserves $\mathscr{O}(H_{d})=[0,1]^{d}$.
The equivariant Hilbert series of $\Delta(B_{d})$ with respect to $\rho$ has been computed by Adams and Reiner [AR23, Example 2.8] and, with this, we can give an alternative expression for the $S_{d}$-equivariant Ehrhart series of $\mathscr{O}(H_{d})$ that was first computed in [Sta11, §9] (see also [Ste94]):
(22) $\operatorname{EE}_{\rho^{*}}({\mathscr{O}{(H_{d})}},{t})=\frac{\sum_{\lambda}\chi_{\lambda}t^{\operatorname{des}(\lambda)}}{(1-t)^{d+1}},$
@@
For $k>0$ let $X_{k}$ be the “radio-tower” poset with $2k$ elements as in Figure 1, where the poset $J(X_{k})$ is also depicted. On $X_{k}$ we consider the action of the group $(S_{2})^{k}$, where the $i$-th factor swaps $a_{i}$ and $b_{i}$. The induced action on $J(X_{i})$ fixes the elements $C_{i}$ and switches $A_{i}$ with $B_{i}$.
[…]
@@
Figure 1. The “radio-tower poset” $X_{k}$ (left-hand side) and its poset of ideals $J(X_{k})$ (right-hand side). We set $A_{i}:=(X_{k})_{\leq a_{i}}$, $B_{i}:=(X_{k})_{\leq b_{i}}$ for all $i\geq 0$, $C_{i}:=A_{i}\cup B_{i}$ for all $i>0$, and $C_{0}:=\emptyset$.
@@
Note that $J(X_{k})$ is ranked, the rank of $C_{i}$ equals $2i$ and the rank of $A_{i}$ and $B_{i}$ is $12i-$ 1. In order to compute the equivariant Ehrhart series of the order polytope associated with $X_{k}$, we consider $S\subseteq\{1,\ldots,2k-1\}$ and we study the $G$-representation on the homology of $J(X_{k})_{S}$. If $S$ contains an even number, then the rank-selected poset contains some $C_{i}$ and is contractible. In particular, its top reduced homology is trivial and affords the zero character. Otherwise, $J(X_{k})_{S}$ is isomorphic to $X_{|S|}$ and the associated order complex is, topologically, an $(|S|-1)$-dimensional sphere that can be realized as the boundary complex of the cross-polytope whose vertices are the standard basis vectors of $\mathbb{R}^{|S|}$ and their negatives. The action of a generator of $G$ corresponds to reflecting the cross-polytope about a coordinate hyperplane, and hence it reverses orientation. This defines an (irreducible) representation of $G$ on the (one-dimensional) top homology of $J(X_{k})_{S}$, isomorphic to the representation in which $(\tau_{1},\ldots,\tau_{k})\in(S_{2})^{k}$ acts on $\mathbb{C}$ by multiplication with $\operatorname{sgn}(\tau_{1})\cdots\operatorname{sgn}(\tau_{k})$. Call $\operatorname{sgn}_{j}$ the sign representation of the $j$-th factor in $(S_{2})^{k}$. Then, $\beta_{S}^{J(X_{k})}=\prod_{i\in S}\operatorname{sgn}_{\frac{i+1}{2}}$ in $\operatorname{R}_{G}$. We obtain the following expression for the equivariant Ehrhart series of $\mathscr{O}(X_{k})$ with respect to the permutation action on the coordinates of $\mathbb{R}^{X}$ (note that, since every element of this group is an involution, this is the “usual” permutation representation):
$\operatorname{EE}_{\rho^{*}}({\mathscr{O}(X_{k})},{t})=\frac{1}{(1-t)^{2k+1}}\sum_{S\subseteq[2k]_{\textrm{odd}}}t^{|S|}\prod_{i\in S}\operatorname{sgn}_{\frac{i+1}{2}}=\frac{\prod_{j=1}^{k}(1+\operatorname{sgn}_{j}t)}{(1-t)^{2k+1}}.$
@@
Example 6.9 is a special case of the following general fact. Following [Bjö95], the join $P\ast Q$ of two posets is the poset on the disjoint union $P\sqcup Q$ with order relation $x\leq y$ if $x,y\in P$ and $x\leq_{P}y$, or $x,y\in Q$ and $x\leq_{Q}y$, or $x\in P,y\in Q$. In particular, the $k$-th radio-tower poset $X_{k}$ is the $k$-fold join of the $2$-element antichain.
@@
Let $X_{1},X_{2}$ be two finite posets and consider the action of a finite group $G$ on the join $X_{1}\ast X_{2}$. Then $G$ must preserve $X_{1}$ and $X_{2}$, and
@@
Following standard notation [Sta82], for a ranked poset $P$ of length $\ell$ and $S\subseteq\{1,\ldots,\ell\}$ we write $P_{S}$ for the poset induced on the set of all elements of $P$ whose rank is in $S$. This implies that $\Delta(P_{S})=\Delta(P)|_{S}$. For brevity we write $\mathbf{J}_{12}:=\Delta(J(X_{1}\ast X_{2})_{S})=\Delta(J(X_{1}\ast X_{2}))|_{S}$ and $\mathbf{J}_{i}:=\Delta(J(X_{i})_{S_{i}})=\Delta(J(X_{i}))|_{S_{i}}$ for $i\in\{1,2\}$. Since joins commute with order complexes (see [Bjö95, §9.4], also for a definition of the join of simplicial complexes), we have

### v3

@@ 4.1. The Cohen–Macaulay approach
Let $\mathcal{O}\in\Sigma/G$. Each $\rho(g)$ permutes the faces of $\Sigma$ belonging to the orbit $\mathcal{O}$, and one can thus define the character $\mathfrak{f}_{\mathcal{O}}^{\Sigma}$ of the associated complex permutation representation. (When $\mathcal{O}$ is the orbit of the empty face of $\Sigma$, we set $\mathfrak{f}_{\mathcal{O}}^{\Sigma}=1$.) Defining $\mathfrak{f}_{i-1}^{\Sigma}$ to be the character of the complex permutation representation on the set of all (i1)(i-1)-dimensional faces of $\Sigma$, it follows that $\mathfrak{f}_{i-1}^{\Sigma}=\sum_{\mathrm{rk}(\mathcal{O})=i}\mathfrak{f}^{\Sigma}_{\mathcal{O}}$.
@@
where we are setting $0^{0}=1$ by convention. Since the coloring $\gamma$ is preserved by the group action and the action permutes the vertices of $\Sigma$, we have that
@@
for some $a^{(i)}_{0},\ldots a^{(i)}_{c}\in\mathbb{K}$. Now define a coloring of $\Sigma$ by assigning to every vertex $v$ the unique color $\gamma(v)\in\{0,\ldots,c\}$ such that $v\in Gw_{\gamma(v)}$. This coloring is obviously preserved by $G$. In order to prove that the action of $G$ is translative, it is enough to prove that the coloring $\gamma$ is proper. If $\Sigma$ is 0-dimensional, this is trivially true. Assume thus that $\Sigma$ has dimension at least $1$ and consider any nonempty edge $\sigma=\{v,w\}\in\Sigma$. By the Kind–Kleinschmidt criterion (Lemma 4.4), the vector subspace $\langle\theta_{0}|_{\sigma},\theta_{1}|_{\sigma},\ldots\theta_{d}|_{\sigma}\rangle$ has dimension $2$. By construction, the coefficient of $x_{v}$ (respectively, $x_{w}$) in $\theta_{i}$ is $a^{(i)}_{\gamma(v)}$ (respectively, $a^{(i)}_{\gamma(w)}$). As a consequence, the $2$-dimensional subspace $\langle\theta_{0}|_{\sigma},\theta_{1}|_{\sigma},\ldots\theta_{d}|_{\sigma}\rangle$ is the image of the $2\times(d+1)$ matrix
@@
The proof of Theorem 4.6 shows that the statement holds also for those finite fields $\mathbb{K}$ containing at least as many distinct elements as the number of vertices of $\Sigma$.
@@
For a simple example where Theorem 4.8 can be used, we refer the reader to Example 3.4. Indeed, the ring $R$ in Example 3.4 is the Stanley–Reisner ring over $\mathbb{C}$ of the 0-dimensional (and hence Cohen–Macaulay) simplicial complex $\Sigma$ having two points as its facets.
@@ 4.2. The colorful approach
Let us write $\mathfrak{f}^{\Sigma}_{S}$ for the character of the associated complex permutation representation. Note that $\mathfrak{f}^{\Sigma}_{S}$ is also the character of the permutation representation of $G$ on the standard basis of the vector space $C_{|S|-1}(\Sigma|_{S},\mathbb{C})$ of (|S|1)(|S|-1)-chains of $\Sigma|_{S}$ constructed with the orientation chosen as in Remark 4.10.
@@
Let $S\subseteq\Gamma$. By [Sta79, Theorem 4.3], the (|S|1)(|S|-1)-dimensional color-selected subcomplex $\Sigma|_{S}$ is Cohen–Macaulay, and hence $\widetilde{H}_{i}(\Sigma|_{S},\mathbb{C})$ vanishes for every i<|S|1i<|S|-1. Then $\mathfrak{h}_{S}^{\Sigma}$ is the character of the $G$-representation on $\widetilde{H}_{|S|-1}(\Sigma|_{S},\mathbb{C})$ from Lemma 4.12. ∎
@@ 5.1. The equivariant Ehrhart series of a lattice polytope
[…]
@@
where $h^{*}(P,\rho;t)$ is a power series in $R_{G}[[t]]$ and not necessarily a polynomial! Note that, when all coefficients of $h^{*}(P,\rho;t)$ are effective, the power series $h^{*}(P,\rho;t)$ is actually a polynomial, as follows from considering its evaluation at $\mathrm{id}_{G}$: see e.g. [EKS24, Lemma 2.8]. It is currently open whether polynomiality and effectiveness of $h^{*}(P,\rho;t)$ are equivalent conditions; this is conjectured to be true in [Sta23, Conjecture 1.2].
@@ 6. Order polytopes
In the setting of Remark 6.5, note that if $S\subseteq\{0,1,\ldots,r\}$ contains at least one of 0 and $r$, then $\Delta({Q})|_{S}$ contains at least one of the cone points $\hat{0}$ and $\hat{1}$ and is hence contractible. This implies that its reduced homology vanishes and, thus, $\mathfrak{h}_{S}^{\Delta(Q)}=\mathfrak{b}_{S}^{\Delta(Q)}=0$.
@@
By Lemma 6.4, the action $\rho^{*}$ on $\mathscr{T}_{X}$ is translative and preserves a balanced coloring, while $\rho$ describes the action on $\Delta(J(X))$ induced by the equivariant isomorphism of abstract simplicial complexes $\phi\colon\underline{\mathscr{T}_{X}}\simeq\Delta(J(X))$. To compute the equivariant Ehrhart series of $\mathscr{O}(X)$ with respect to $\rho^{*}$, by Theorem 5.2 it is enough to compute the equivariant Hilbert series of the abstract simplicial complex $\underline{\mathscr{T}_{X}}$ with respect to $\rho^{*}$. By the equivariant isomorphism of abstract simplicial complexes $\phi\colon\underline{\mathscr{T}_{X}}\simeq\Delta(J(X))$ from Lemma 6.4, this is in turn equal to the equivariant Hilbert series of the abstract simplicial complex $\Delta(J(X))$ with respect to $\rho$. The poset $J(X)$ is bounded with $\hat{0}$$\emptyset$ and $\hat{1}=X$, and it is ranked by cardinality of its elements. By Remark 6.6, we can restrict to color sets $S\subseteq[d-1]$. The claim about the equivariant flag $h$-vector refining the equivariant $h$-vector follows from Proposition 4.14.(ii). ∎
Following standard notation [Sta82], for a ranked poset $P$ of length $\ell$ and $S\subseteq\{1,\ldots,\ell\}$ we write $P_{S}$ for the poset induced on the set of all elements of $P$ whose rank is in $S$. This implies that $\Delta(P_{S})=\Delta(P)|_{S}$.
@@
Let $\mathsf{Anti}_{d}$ denote the poset consisting of a single antichain with $d$ elements. Then $\mathscr{O}(\mathsf{Anti}_{d})$ is the unit cube $0,1]^{d}$. Consider the group $G=S_{d}$, the symmetric group on $d$ elements, acting on $\mathsf{Anti}_{d}$. The poset $J(\mathsf{Anti}_{d})$ is the boolean poset $B_{d}$ of all subsets of $[d]$, and the induced action of $S_{d}$ is the natural action permuting elements of $[d]$. As in Theorem 6.7, let $\rho$ denote the corresponding action on the Stanley-Reisner ring of $\Delta(J(\mathsf{Anti}_{d}))=\Delta(B_{d})$ and let $\rho^{*}$ denote the representation of $S_{d}$ permuting coordinates in $\mathbb{R}^{\mathsf{Anti}_{d}}$. Then $\rho^{*}$ preserves $\mathscr{O}(\mathsf{Anti}_{d})=[0,1]^{d}$.
The equivariant Hilbert series of $\Delta(B_{d})$ with respect to $\rho$ has been computed by Adams and Reiner [AR23, Example 2.8] and, with this, we can give an alternative expression for the $S_{d}$-equivariant Ehrhart series of $\mathscr{O}(\mathsf{Anti}_{d})$ that was first computed in [Sta11, §9] (see also [Ste94]):
(22) $\operatorname{EE}_{\rho^{*}}({\mathscr{O}{(\mathsf{Anti}_{d})}},{t})=\frac{\sum_{\lambda}\chi_{\lambda}t^{\operatorname{des}(\lambda)}}{(1-t)^{d+1}},$
@@
For $k>0$ let $\mathsf{Rad}_{k}$ be the “radio-tower” poset with $2k$ elements as in Figure 1, where the poset $J(\mathsf{Rad}_{k})$ is also depicted. We set $A_{i}:=(\mathsf{Rad}_{k})_{\leq a_{i}}$, $B_{i}:=(\mathsf{Rad}_{k})_{\leq b_{i}}$ for all $i\geq 0$, $C_{i}:=A_{i}\cup B_{i}$ for all $i>0$, and $C_{0}:=\emptyset$. On $\mathsf{Rad}_{k}$ we consider the action of the group $(S_{2})^{k}$, where the $i$-th factor swaps $a_{i}$ and $b_{i}$. The induced action on $J(\mathsf{Rad}_{i})$ fixes the elements $C_{i}$ and switches $A_{i}$ with $B_{i}$.
$a_{1}$$a_{2}$$a_{3}$$a_{k}$$b_{1}$$b_{2}$$b_{3}$$b_{k}$$A_{1}$$A_{2}$$A_{3}$$B_{1}$$B_{2}$$B_{3}$$C_{1}$$C_{2}$$C_{k}$$C_{0}=\emptyset$
@@
Figure 1. The “radio-tower poset” $\mathsf{Rad}_{k}$ (left-hand side) and its poset of ideals $J(\mathsf{Rad}_{k})$ (right-hand side).
@@
Note that $J(\mathsf{Rad}_{k})$ is ranked, the rank of $C_{i}$ equals $2i$ and the rank of $A_{i}$ and $B_{i}$ is 2i12i-1. In order to compute the equivariant Ehrhart series of the order polytope associated with $\mathsf{Rad}_{k}$, we consider $S\subseteq\{1,\ldots,2k-1\}$ and we study the $G$-representation on the homology of $J(\mathsf{Rad}_{k})_{S}$. If $S$ contains an even number, then $J(\mathsf{Rad}_{k})_{S}$ contains some $C_{i}$ and is contractible. In particular, its top reduced homology is trivial and affords the zero character. Otherwise, $J(\mathsf{Rad}_{k})_{S}$ is isomorphic to $\mathsf{Rad}_{|S|}$ and the associated order complex is, topologically, an (|S|1)(|S|-1)-dimensional sphere that can be realized as the boundary complex of the cross-polytope whose vertices are the standard basis vectors of $\mathbb{R}^{|S|}$ and their negatives. The action of a generator of $G$ corresponds to reflecting the cross-polytope about a coordinate hyperplane, and hence it reverses orientation. This defines an (irreducible) representation of $G$ on the (one-dimensional) top homology of $J(\mathsf{Rad}_{k})_{S}$, isomorphic to the representation in which $(\tau_{1},\ldots,\tau_{k})\in(S_{2})^{k}$ acts on $\mathbb{C}$ by multiplication with $\operatorname{sgn}(\tau_{1})\cdots\operatorname{sgn}(\tau_{k})$. Call $\operatorname{sgn}_{j}$ the sign representation of the $j$-th factor in $(S_{2})^{k}$. Then, $\beta_{S}^{J(\mathsf{Rad}_{k})}=\prod_{i\in S}\operatorname{sgn}_{\frac{i+1}{2}}$ in $\operatorname{R}_{G}$. We obtain the following expression for the equivariant Ehrhart series of $\mathscr{O}(\mathsf{Rad}_{k})$ with respect to the permutation action on the coordinates of $\mathbb{R}^{\mathsf{Rad}_{k}}$ (note that, since every element of this group is an involution, this is the “usual” permutation representation):
$\operatorname{EE}_{\rho^{*}}({\mathscr{O}(\mathsf{Rad}_{k})},{t})=\frac{1}{(1-t)^{2k+1}}\sum_{S\subseteq[2k]_{\textrm{odd}}}t^{|S|}\prod_{i\in S}\operatorname{sgn}_{\frac{i+1}{2}}=\frac{\prod_{j=1}^{k}(1+\operatorname{sgn}_{j}t)}{(1-t)^{2k+1}}.$
@@
Example 6.9 is a special case of the following general fact. Following [Bjö95, §9.4], the ordinal sum $P\ast Q$ of two posets is the poset on the disjoint union $P\sqcup Q$ with order relation $x\leq y$ if $x,y\in P$ and $x\leq_{P}y$, or $x,y\in Q$ and $x\leq_{Q}y$, or $x\in P,y\in Q$. In particular, the $k$-th radio-tower poset $\mathsf{Rad}_{k}$ is the $k$-fold ordinal sum of the $2$-element antichain. If $\Sigma_{1}$ and $\Sigma_{2}$ are two abstract simplicial complexes with no vertices in common, we denote by $\Sigma_{1}\ast\Sigma_{2}$ their join, i.e., the abstract simplicial complex whose faces are the collections $\sigma_{1}\cup\sigma_{2}$ with $\sigma_{1}\in\Sigma_{1}$ and $\sigma_{2}\in\Sigma_{2}$.
@@
Let $X_{1},X_{2}$ be two finite posets and consider the action of a finite group $G$ on the ordinal sum $X_{1}\ast X_{2}$. Then $G$ must preserve $X_{1}$ and $X_{2}$, and
@@
For brevity we write $\mathbf{J}_{12}:=\Delta(J(X_{1}\ast X_{2})_{S})=\Delta(J(X_{1}\ast X_{2}))|_{S}$ and $\mathbf{J}_{i}:=\Delta(J(X_{i})_{S_{i}})=\Delta(J(X_{i}))|_{S_{i}}$ for $i\in\{1,2\}$. Since the order complex of an ordinal sum of posets is the join of the order complexes of those posets (see [Bjö95, §9.4]), we have